Exercise 4-26 (Algorithmic) (LO. 4)

Determine the taxable amount of social security benefits for the following situations. If required, round your answers to the nearest dollar. If an amount is zero, enter "0".

a. Erwin and Eleanor are married and file a joint tax return. They have adjusted gross income of $39,400, no tax-exempt interest, and $13,790 of Social Security benefits. As a result, $___________ of the Social Security benefits are taxable.

My thoughts were that the answer is whichever is the lesser of  50% of Social Security benefits ($6895) or AGI ($39400) - $32000 = 7400*50% = $3700. However, both these calculations are incorrect. Please any assistance would be greatly appreciated, Thank You!!!!
Answer #1

They shall include $6,895 of the Social Security benefit in gross income. This works out as the lesser of the following:

>> 0.50 ($13,790 Social Security benefits) = $6,895

>>0.50 [$39,400+ 0.50($13,790) - 32,000]

= 0.50 ($14,295)= $7,148

IN CASE OF ANY DOUBTS FEEL FREE TO COMMENT....PLEASE GIVE ME A ??
Answer #2

$633*12*20.8= $157,997

76000/157997= 0.4810

7600*0.4810=3656  Exclusion amount

7600-3656=3944 included